Three resistors A,B , and C are connected in series to a 117 volt source . If RA= 64 ohms , and EB = 40 volts when the current is 0.5 amp , calculate the resistors RB and RC


1
Expert's answer
2021-09-24T02:14:50-0400

"-117 + 64(0.5) + 40 + R_c(0.5) = 0\\
R_c = \frac{45}{0.5} = 90ohms\\
R_b = \frac{40}{0.5} = 80 ohms\\"